OpenStudy (anonymous):

please help me i dont understand this ? :( The equation on a line is 2x - 2y = 6 what is the slope of the line? A.-3 B.-1 C. 1/3 D. 1

OpenStudy (skullpatrol):

Failure to follow the guidelines may lead to your suspension or banning from the site. If you see any violations, please report them by using a report abuse link. OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

with an algebraic manipulations make this equation look like y=mx+b then slope will be m, or just diferentiate w.r.t. x and thats it.

OpenStudy (davidusa):

@skullpatrol that is so meaan

OpenStudy (anonymous):

@DavidUsa lol ikr

OpenStudy (anonymous):

@le0n what is w.r.t ? :/ n sorry for asking so many questions im just so overwhelmed..

OpenStudy (anonymous):

with respect to

OpenStudy (anonymous):

its an abriviation

OpenStudy (ybarrap):

Make the coefficient in front of the y equal to "1". What do we need to divide the whole thing by to make this happen? Once you do this, move the variable x to the other side of the equation. Look at the coefficient in front of the x variable. That's your slope.

OpenStudy (ybarrap):

Hint: Divide everything by -2. Now the coefficient in front of the y variable is "1". Divide all the other numbers by -2 also.

OpenStudy (ybarrap):

We have: \(\Large 2x - 2y = 6\): Divide both sides by -2 \( \Large {2x - 2y\over -2} = \dfrac 6 {-2} \) \( \Large \dfrac{2x}{-2} +y = \dfrac 6 {-2} \): Now we have y with a 1 coefficient Now what happens to the coeficient in front of the x variable? That is, what's its value Move that variable to the other side of the equation. Then just look what happened to the x coefficient. That's your slope.

OpenStudy (ybarrap):

Does this make sense?

OpenStudy (ybarrap):

\(\Large y = \dfrac 6 {-2}+\dfrac{2x}{2} \): Here's your answer, just simplify the coefficient in front of the x and you're done.